OpenStudy (anonymous):

how do you solve this? 7/3 - 1/x-1= -3/3x-3 these are fractions

OpenStudy (lynfran):

\[\frac{ 7 }{ 3 }-\frac{ 1 }{ x-1 }=\frac{ -3 }{ 3x-3 }\]..is this it?

OpenStudy (anonymous):

yeah ^^

OpenStudy (lynfran):

\[\frac{ 7x-7-3 }{ 3(x-1) }=\frac{ -3 }{ 3x-3 }\]\[\frac{ 7x-10 }{ 3x-3 }=\frac{ -3 }{ 3x-3 }\].....no solutions to this problem... http://www.mathpapa.com/algebra-calculator.html

OpenStudy (lynfran):

@alysha_32